# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ Abstract
Keywords: Analysis of covariance; causal inference; cluster-robust variance estimator; covariate adjustment; efficiency comparison; heteroskedasticity-consistent variance estimator
@@ 1 Introduction
We study a class of randomized experiments with clusters as the unit of randomization, with the treatment rolled out over multiple periods. At baseline, each cluster is randomized to a specific treatment adoption time, and clusters that have already begun treatment will continue to receive treatment until the end of the study. We refer to this design as the staggered rollout cluster randomized experiment (SR-CREs), which is attractive in settings where simultaneous randomization in a single period is not logistically feasible. SR-CREs have also been referred to as cluster randomized difference-in-differences (DID) in social science (Athey and Imbens, 2022) and stepped wedge cluster randomized trials in medicine (Hussey and Hughes, 2007), and have become increasingly common.
@@
Previous staggered rollout design literature primarily focused on single-level data, e.g., observational DID for panel data where the treatment assignment and outcome measurements occur at the same level (Callaway and Sant’Anna, 2021; Sun and Abraham, 2021) and staggered rollout individual randomized experiments (SR-IREs, Athey and Imbens, 2022; Bojinov et al., 2021; Roth and Sant’Anna, 2023). Differentiating the level of assignment from the level of outcome measurements, there is an emerging literature on design-based inference for CREs without staggered rollout. For example, Schochet et al. (2022) proposed design-based estimators for blocked CREs. Su and Ding (2021) developed a unified theory for design-based model-assisted inference in CREs under parallel-arm assignments. Differentiating the level of treatment assignment from the level of outcome measurements under staggered rollout, Schochet (2022) studied the statistical power of DID and comparative interrupted time series estimators for panel data from observational studies. Chen and Li (2024) proposed a class of weighted average treatment effect estimands and studied the design-based properties of analysis of covariance (ANCOVA) estimators for stepped wedge CREs. However, these developments assumed away anticipation (Malani and Reif, 2015, a psychological phenomenon where clusters bring the expected future into the present) and dynamic causal effects (Sun and Abraham, 2021, treatment effects evolve over length of exposure), which cannot be easily ruled out a priori in practice.
@@
Focusing on the central question of treatment effect evaluation, this article addresses regression estimators for SR-CREs from a design-based perspective, where the potential outcomes and covariates are assumed to be fixed, and the randomness arises solely from the randomized treatment adoption times. We propose a class of weighted average treatment effect estimands that account for the unique design features of an SR-CRE. Specifically, we relax the no-anticipation assumption (Athey and Imbens, 2022; Chen and Li, 2024; Roth and Sant’Anna, 2023) and allow for dynamic causal effects (Sun and Abraham, 2021; Wang et al., 2024), such that the treatment adoption time is an intrinsic component of the estimands. Similar to Su and Ding (2021) and Chen and Li (2024), we also explicitly address the cluster-period size as an integral part of the estimands and generalize the dynamic DID estimands (Sun and Abraham, 2021; Callaway and Sant’Anna, 2021) to address differential levels of assignment and data measurement as well as non-ignorable cluster-period sizes. Based on the proposed estimands, we can then define the dynamic causal effects, which compare the summary of potential outcomes between any pair of treatment adoption times in specific calendar periods to address different scientific questions. To our knowledge, this is the first effort to simultaneously address anticipation, dynamic causal effects, and non-ignorable cluster-period sizes in estimations and definitions for SR-CREs.
@@ 2.1 Identification assumptions
where $\bm{A}=(A_{1},\ldots,A_{I})^{T}$ and $\bm{a}=(a_{1},\ldots,a_{I})^{T}$ with $\sum_{i=1}^{I}\mathbbm{1}(A_{i}=a)=I(a)$.
This assumption states that the treatment adoption times are random. It is common in the staggered rollout design literature, e.g., Athey and Imbens (2022, Assumption 1), Roth and Sant’Anna (2023, Assumption 1), and Chen and Li (2024, Assumption 3), and is usually held by design.
@@ 2.2 Dynamic causal effect estimands
$\tau_{j}(a,a^{\prime})=\sum_{i=1}^{I}\sum_{k=1}^{N_{ij}}\pi_{ijk}\{Y_{ijk}(a)-Y_{ijk}(a^{\prime})\};$ (2)
and using the cluster-period averages as
@@
The definition of $\tau_{j}(a,a^{\prime})$ in (1) implies that $\tau_{j}(a^{\prime},a)=\overline{Y}_{\cdot j\cdot}(a^{\prime})-\overline{Y}_{\cdot j\cdot}(a)=-\tau_{j}(a,a^{\prime})$ and $\tau_{j}(a,a)=0$. Therefore, it is sufficient to focus on $\tau_{j}(a,a^{\prime})$’s with $a<a^{\prime}$. Depending on the comparative relationship between period $j$ and treatment adoption times $a$ and $a^{\prime}$, the DWATE estimand $\tau_{j}(a,a^{\prime})$ has the following causal interpretations summarized in Table 1. Causal interpretations of $\tau_{j}(a,a^{\prime})$ are straightforward for $j<a<a^{\prime}$ and $a<a^{\prime}\leq j$, as they capture fine-resolution anticipation and duration effects, respectively. Of note, there are no anticipation effects when $j=J$ or $a=1$, and naturally, the anticipation effects are assumed to exist only for eventually treated clusters. For $a\leq j<a^{\prime}$, causal interpretations of $\tau_{j}(a,a^{\prime})$ could be more subtle for a general $a^{\prime}$ because of the anticipation component brought by the relationship $j<a^{\prime}$. Nevertheless, we provide a discussion on the general $\tau_{j}(a,a^{\prime})$ for theoretical interests and completeness of the subsequent development. To focus ideas, the causal interpretations in this scenario can be more streamlined by setting $a^{\prime}=\infty$, allowing us to draw comparisons to the average potential outcomes had the clusters never been treated (an imaginary counterfactual control condition). In this case, we can define $\mathrm{WATE}_{j}(a)=\tau_{j}(a,\infty)$ for $a\leq j$ as the weighted average treatment effect (WATE) of treatment adoption time $a$ measured for the population in calendar period $j$. The WATE is analogous to but differs from the average treatment effect on the treated under non-random treatment adoption times (Callaway and Sant’Anna, 2021; Sun and Abraham, 2021), as the latter is defined on the treated group. Furthermore, when setting $a^{\prime}=\infty$, we can also define $\mathrm{AWATE}_{j}(a)=\tau_{j}(a,\infty)$ for $j<a$ as the anticipated WATE (AWATE) of treatment adoption time $a$ measured for the population in calendar period $j$ (Abbring and Van Den Berg, 2003; Malani and Reif, 2015).
Remark 1
The estimand $\tau_{j}(a,a^{\prime})$ and its causal interpretations can be simplified if certain auxiliary assumptions are adopted in addition to Assumptions 1 and 2. For example, if we assume no anticipation (Roth and Sant’Anna, 2023; Schochet, 2022, Assumption 2), i.e., for all $i$ and $k$, $Y_{ijk}(a)=Y_{ijk}(a^{\prime})$ for all $j<\min(a,a^{\prime})$, then the anticipation effects are excluded, because $\tau_{j}(a,a^{\prime})=0$ for $j<a<a^{\prime}$ and $\tau_{j}(a,j+1)=\cdots=\tau_{j}(a,J)=\tau_{j}(a,\infty)$ for $a\leq j$. If we further consider the invariance to history assumption (Athey and Imbens, 2022, Assumption 3), i.e., for all $i$ and $k$, $Y_{ijk}(a)=Y_{ijk}(1)$ for all $a\leq j$, then, by Lemma 2 of Athey and Imbens (2022), the duration effects are also eliminated, because, for all $i$ and $k$, $Y_{ijk}(a)=Y_{ijk}(1)$ for all $a\leq j$ and $Y_{ijk}(a)=Y_{ijk}(\infty)$ for all $a>j$. Therefore, $\tau_{j}(a,a^{\prime})=\tau_{j}(1,\infty)$ for $a\leq j<a^{\prime}$ and $\tau_{j}(a,a^{\prime})=0$ otherwise; this is the assumption taken in the majority of the existing literature for SR-CREs to simplify the treatment effect estimands (Chen and Li, 2024, Assumption 2). Alternatively, the less restrictive limited treatment anticipation assumption (Callaway and Sant’Anna, 2021, Assumption 3) has also been employed in the staggered rollout design literature, where an anticipation horizon $\delta\geq 0$ is assumed such that $Y_{ijk}(a)=Y_{ijk}(a^{\prime})$ for all $j<\min(a,a^{\prime})-\delta$ with $\tau_{j}(a,a^{\prime})=0$ for $j<\min(a,a^{\prime})-\delta$. The limited treatment anticipation assumption reduces to the no anticipation assumption when $\delta=0$.
@@
which is a weighted sum of $\tau_{j}(a,a^{\prime})$ with $b_{\theta,j}(a,a^{\prime})\in\mathbb{R}$ being user-defined weights. For example, with $\tau_{j}(a,\infty)=\mathrm{WATE}_{j}(a)$ ($a\leq j$), we can define a grand average estimand over adoption and calendar times as $\mathrm{OWTE}^{sim}=\sum_{j=1}^{J}\sum_{a:a\leq j}w_{\cdot j\cdot}I(a)\mathrm{WATE}_{j}(a)/\allowbreak\sum_{j=1}^{J}\sum_{a:a\leq j}w_{\cdot j\cdot}I(a)$. Similarly, Using $\tau_{j}(a,\infty)=\mathrm{AWATE}_{j}(a)$ ($j<a$) as building blocks, a grand average estimand with a focus on anticipation effects is given by $\mathrm{OAWTE}^{sim}=\sum_{j=1}^{J-1}\sum_{a:j<a\leq J}\allowbreak w_{\cdot j\cdot}I(a)\mathrm{AWATE}_{j}(a)/\sum_{j=1}^{J-1}\sum_{a:j<a\leq J}w_{\cdot j\cdot}I(a)$. In Section S1 of the Online Supplement, we provide four classes of summary estimands to address different scientific questions.
@@ 3.1 Preliminaries
Let $\bm{\xi}_{i\cdot}(a)=(\xi_{i1\cdot}(a),\ldots,\xi_{iJ\cdot}(a))^{T}$ for $a\in\mathcal{A}$ and $i=1,\ldots,I$ denote general cluster-level potential outcomes with zero averages over clusters. Define the individual centered potential outcome $\epsilon_{ijk}(a)=Y_{ijk}(a)-\overline{Y}_{\cdot j\cdot}(a)$, and then examples of $\xi_{ij\cdot}(a)$ include the cluster-period average centered potential outcome, $\overline{\epsilon}_{ij\cdot}(a)=\overline{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)$, and the scaled cluster-period total, $\widetilde{\epsilon}_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)$. We then define the following matrices to streamline the exposition of the theoretical properties of the estimators, where
$\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})=\sum_{i=1}^{I}\left\{\frac{\bm{\xi}_{i\cdot}(a)\bm{\xi}_{i\cdot}(a)^{T}}{I(a)}+\frac{\bm{\xi}_{i\cdot}(a^{\prime})\bm{\xi}_{i\cdot}(a^{\prime})^{T}}{I(a^{\prime})}\right\}$
and $\bm{V}(\bm{\xi}_{i\cdot})(a,a^{\prime})=\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})-I^{-1}\sum_{i=1}^{I}\left\{\bm{\xi}_{i\cdot}(a)-\bm{\xi}_{i\cdot}(a^{\prime})\right\}\left\{\bm{\xi}_{i\cdot}(a)-\bm{\xi}_{i\cdot}(a^{\prime})\right\}^{T}$. For two treatment adoption times $a$ and $a^{\prime}$, $\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ and $\bm{V}(\xi_{i\cdot})(a,a^{\prime})$ are covariance matrices of the difference in means of $\{\bm{\xi}_{i\cdot}(a),\bm{\xi}_{i\cdot}(a^{\prime})\}_{i=1}^{I}$ multiplied by (I1)(I-1) in completely randomized experiments with and without the constant treatment effect assumption. In a more granular fashion, examining the $(j,j)$ entry of $\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ and $\bm{V}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ gives us
@@ 3.2.1 Without covariates
Let $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ denote the covariance matrix of $\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})=(\widehat{\tau}_{1,\mathrm{I}}(a,a^{\prime}),\ldots,\widehat{\tau}_{J,\mathrm{I}}(a,a^{\prime}))^{T}$ and $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ denote the marginal variance of $\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})$, where $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ is the $(j,j)$ entry of $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$. Both $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ and $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ can be estimated by the CR variance estimator (Liang and Zeger, 1986), denoted by $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ and $\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$, respectively. For a WLS fit of (6), let $\bm{D}_{i,\mathrm{I}}$ and $\bm{\Pi}_{i,\mathrm{I}}$ respectively denote the design matrix and diagonal weight matrix for cluster $i$. We first obtain the following CR variance estimator for the vector of the regression coefficients,
$\bm{E}_{\mathrm{I}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{T}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)^{-1},$
@@
where $\widehat{\bm{e}}_{i,\mathrm{I}}$ is the $N_{i\cdot}$-dimensional vector of residuals from the WLS fit of (6), with $N_{i\cdot}=\sum_{j=1}^{J}N_{ij}$, $\bm{D}_{i,\mathrm{I}}$ is of dimensions $N_{i\cdot}\times J(J+1)$ with $\bm{D}_{i,\mathrm{I}}=(\bm{D}_{i1,\mathrm{I}}^{T}\cdots\bm{D}_{iJ,\mathrm{I}}^{T})^{T}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}$ being $\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ijk}^{T}$, and the $j$-th diagonal block of $\bm{\Pi}_{i,\mathrm{I}}$ is $\bm{\Pi}_{ij,\mathrm{I}}=\mathrm{diag}(\pi_{ij1},\ldots,\pi_{ijN_{ij}})$. Here, $\bm{G}_{i}(\mathcal{A})=(G_{i}(1),\ldots,G_{i}(J),G_{i}(\infty))^{T}$, $\overrightarrow{\mathbbm{1}}_{ijk}=(\mathbbm{1}_{ijk,1}\ldots,\mathbbm{1}_{ijk,J})^{T}$ with $\mathbbm{1}_{ijk,j^{\prime}}=1$ if $j=j^{\prime}$ and $\mathbbm{1}_{ijk,j^{\prime}}=0$ otherwise for $j^{\prime}=1,\ldots,J$, and $\otimes$’ denotes the Kronecker product such that $\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ijk}^{T}=(G_{i}(1)\mathbbm{1}_{ijk}^{T},\ldots,G_{i}(J)\mathbbm{1}_{ijk}^{T},\allowbreak G_{i}(\infty)\mathbbm{1}_{ijk}^{T})$. We can write $\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})=\bm{Q}(a,a^{\prime})\widehat{\bm{\beta}}_{\mathrm{I}}$, where $\widehat{\bm{\beta}}_{\mathrm{I}}=(\widehat{\bm{\beta}}_{\mathrm{I}}(1)^{T},\ldots,\widehat{\bm{\beta}}_{\mathrm{I}}(J)^{T},\widehat{\bm{\beta}}_{\mathrm{I}}(\infty)^{T})^{T}$, with $\widehat{\bm{\beta}}_{\mathrm{I}}(a)=(\widehat{\beta}_{1,\mathrm{I}}(a),\ldots,\widehat{\beta}_{J,\mathrm{I}}(a))^{T}$ for $a\in\mathcal{A}$, and $\bm{Q}(a,a^{\prime})$ is a $J\times J(J+1)$-dimensional known coefficient matrix, e.g.,

### v3

@@ Abstract
Keywords: Causal inference; cluster-robust variance estimator; covariate adjustment; randomized experiments; staggered rollout
@@ 1 Introduction
We study a class of randomized experiments in which clusters are the unit of randomization, with the treatment rolled out over multiple periods. At baseline, each cluster is randomized to a specific treatment adoption time, and clusters that have already begun treatment will continue to receive treatment until the end of the study. We refer to this design as the staggered rollout cluster-randomized experiment (SR-CRE), which is attractive in settings where simultaneous randomization in a single period is not logistically feasible. SR-CREs have also been referred to as cluster-randomized difference-in-differences (DID) in the social sciences (Athey and Imbens, 2022) and stepped-wedge cluster-randomized trials in medicine (Hussey and Hughes, 2007), and have become increasingly common.
@@
Previous staggered rollout design literature has primarily focused on single-level data, e.g., observational DID for panel data, where the treatment assignment and outcome measurements occur at the same level (Callaway and Sant’Anna, 2021; Sun and Abraham, 2021), and staggered rollout individual randomized experiments (SR-IREs, Athey and Imbens, 2022; Bojinov et al., 2021; Roth and Sant’Anna, 2023). Differentiating the level of assignment from the level of outcome measurements, there is an emerging literature on design-based inference for CREs without staggered rollout. For example, Schochet et al. (2022) proposed design-based estimators for blocked CREs. Su and Ding (2021) developed a unified theory for design-based model-assisted inference in CREs under parallel-arm assignments. Differentiating the level of treatment assignment from the level of outcome measurements under staggered rollout, Schochet (2022) studied the statistical power of DID and comparative interrupted time series estimators for panel data from observational studies. Chen and Li (2024) proposed a class of weighted average treatment effect estimands and studied the design-based properties of analysis of covariance (ANCOVA) estimators for stepped wedge CREs. However, these developments assumed away anticipation (Malani and Reif, 2015, a psychological phenomenon where clusters bring the expected future into the present) and dynamic causal effects (Sun and Abraham, 2021, treatment effects evolve over length of exposure), which cannot be easily ruled out a priori in practice.
@@
Focusing on the central question of treatment effect evaluation, this article addresses regression estimators for SR-CREs from a design-based perspective, where the potential outcomes and covariates are assumed fixed, and the randomness arises solely from the randomized treatment adoption times. We propose a class of weighted average treatment effect estimands that account for the unique design features of an SR-CRE. Specifically, we largely relax the no-anticipation assumption (Athey and Imbens, 2022; Chen and Li, 2024; Roth and Sant’Anna, 2023) and allow for dynamic causal effects (Sun and Abraham, 2021; Wang et al., 2024), such that the treatment adoption time is an intrinsic component of the estimands. Similar to Su and Ding (2021) and Chen and Li (2024), we also explicitly address the cluster-period size as an integral part of the estimands and generalize the dynamic DID estimands (Sun and Abraham, 2021; Callaway and Sant’Anna, 2021) to address differential levels of assignment and data measurement as well as non-ignorable cluster-period sizes. Based on the proposed estimands, we can then define the dynamic causal effects, which compare the summary of potential outcomes between any pair of treatment adoption times in specific calendar periods to address different scientific questions. To our knowledge, this is the first effort to simultaneously address anticipation, dynamic causal effects, and non-ignorable cluster-period sizes in estimations and definitions for SR-CREs.
@@ 2.1 Identification assumptions
where $\bm{A}=(A_{1},\ldots,A_{I})^{\top}$ and $\bm{a}=(a_{1},\ldots,a_{I})^{\top}$ with $\sum_{i=1}^{I}\mathbbm{1}(A_{i}=a)=I(a)$.
This assumption states that the treatment adoption times are random. It is a common assumption in the staggered rollout design literature, e.g., Athey and Imbens (2022, Assumption 1), Roth and Sant’Anna (2023, Assumption 1), and Chen and Li (2024, Assumption 3), and is typically held by design.
@@ 2.2 Dynamic causal effect estimands
$\tau_{j}(a,a^{\prime})=\sum_{i=1}^{I}\sum_{k=1}^{N_{ij}}\pi_{ijk}\{Y_{ijk}(a)-Y_{ijk}(a^{\prime})\},$ (2)
with $\overline{Y}_{\cdot j\cdot}(a)=\sum_{i=1}^{I}\sum_{k=1}^{N_{ij}}\pi_{ijk}Y_{ijk}(a)$; and using the cluster-period averages as
@@
The definition of $\tau_{j}(a,a^{\prime})$ in (1) implies that $\tau_{j}(a^{\prime},a)=\overline{Y}_{\cdot j\cdot}(a^{\prime})-\overline{Y}_{\cdot j\cdot}(a)=-\tau_{j}(a,a^{\prime})$ and $\tau_{j}(a,a)=0$. Therefore, it is sufficient to focus on $\tau_{j}(a,a^{\prime})$’s with $a<a^{\prime}$. Depending on the comparative relationship between period $j$ and treatment adoption times $a$ and $a^{\prime}$, the DWATE estimand $\tau_{j}(a,a^{\prime})$ has the following causal interpretations summarized in Table 1. Causal interpretations of $\tau_{j}(a,a^{\prime})$ are straightforward for $j<a<a^{\prime}$ and $a<a^{\prime}\leq j$, as they capture fine-resolution anticipation and duration effects, respectively. Of note, there are no anticipation effects when $j=J$ or $a=1$, and naturally, the anticipation effects are assumed to exist only for eventually treated clusters. For $a\leq j<a^{\prime}$, causal interpretations of $\tau_{j}(a,a^{\prime})$ could be more subtle for a general $a^{\prime}$ because of the anticipation component brought by the relationship $j<a^{\prime}$. Nevertheless, we provide a discussion on the general $\tau_{j}(a,a^{\prime})$ for theoretical interests and completeness of the subsequent development. To focus ideas, the causal interpretations in this scenario can be more streamlined by setting $a^{\prime}=\infty$, allowing us to draw comparisons to the average potential outcomes had the clusters never been treated (an imaginary counterfactual control condition). In this case, we can define $\mathrm{WATE}_{j}(a)=\tau_{j}(a,\infty)$ for $a\leq j$ as the weighted average treatment effect (WATE) of treatment adoption time $a$ measured for the population in calendar period $j$. The WATE is analogous to but differs from the average treatment effect on the treated under non-random treatment adoption times (Callaway and Sant’Anna, 2021; Sun and Abraham, 2021), as the latter is defined for the treated group. Furthermore, when setting $a^{\prime}=\infty$, we can also define $\mathrm{AWATE}_{j}(a)=\tau_{j}(a,\infty)$ for $j<a$ as the anticipated WATE (AWATE) of treatment adoption time $a$ measured for the population in calendar period $j$ (Abbring and Van Den Berg, 2003; Malani and Reif, 2015). Additional special cases of our estimands can be found in Remark S1 in Section S1 of the Online Supplement.
@@
which is a weighted sum of $\tau_{j}(a,a^{\prime})$ with $b_{\theta,j}(a,a^{\prime})\in\mathbb{R}$ being user-defined weights. For example, with $\tau_{j}(a,\infty)=\mathrm{WATE}_{j}(a)$ ($a\leq j$), we can define a grand average estimand over adoption and calendar times as $\mathrm{OWTE}^{sim}=\sum_{j=1}^{J}\sum_{a:a\leq j}w_{\cdot j\cdot}I(a)\mathrm{WATE}_{j}(a)/\allowbreak\sum_{j=1}^{J}\sum_{a:a\leq j}w_{\cdot j\cdot}I(a)$. Similarly, Using $\tau_{j}(a,\infty)=\mathrm{AWATE}_{j}(a)$ ($j<a$) as building blocks, a grand average estimand with a focus on anticipation effects is given by $\mathrm{OAWTE}^{sim}=\sum_{j=1}^{J-1}\sum_{a:j<a\leq J}\allowbreak w_{\cdot j\cdot}I(a)\mathrm{AWATE}_{j}(a)/\sum_{j=1}^{J-1}\sum_{a:j<a\leq J}w_{\cdot j\cdot}I(a)$. We provide four classes of summary estimands to address different scientific questions in Section S1 of the Online Supplement.
@@ 3.1 Preliminaries
Let $\bm{\xi}_{i\cdot}(a)=(\xi_{i1\cdot}(a),\ldots,\xi_{iJ\cdot}(a))^{\top}$ for $a\in\mathcal{A}$ and $i=1,\ldots,I$ denote general cluster-level potential outcomes with zero averages over clusters. Define the individual centered potential outcome $\epsilon_{ijk}(a)=Y_{ijk}(a)-\overline{Y}_{\cdot j\cdot}(a)$, and then examples of $\xi_{ij\cdot}(a)$ include the cluster-period average centered potential outcome, $\overline{\epsilon}_{ij\cdot}(a)=\overline{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)$, and the scaled cluster-period total, $\widetilde{\epsilon}_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)$. We then define the following matrices to streamline the exposition of the theoretical properties of the estimators, where
$\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})=\sum_{i=1}^{I}\left\{\frac{\bm{\xi}_{i\cdot}(a)\bm{\xi}_{i\cdot}(a)^{\top}}{I(a)}+\frac{\bm{\xi}_{i\cdot}(a^{\prime})\bm{\xi}_{i\cdot}(a^{\prime})^{\top}}{I(a^{\prime})}\right\}$
and $\bm{V}(\bm{\xi}_{i\cdot})(a,a^{\prime})=\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})-I^{-1}\sum_{i=1}^{I}\left\{\bm{\xi}_{i\cdot}(a)-\bm{\xi}_{i\cdot}(a^{\prime})\right\}\left\{\bm{\xi}_{i\cdot}(a)-\bm{\xi}_{i\cdot}(a^{\prime})\right\}^{\top}$. For two treatment adoption times $a$ and $a^{\prime}$, $\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ and $\bm{V}(\xi_{i\cdot})(a,a^{\prime})$ are covariance matrices of the difference in means of $\{\bm{\xi}_{i\cdot}(a),\bm{\xi}_{i\cdot}(a^{\prime})\}_{i=1}^{I}$ multiplied by (I1)(I-1) in completely randomized experiments with and without the constant treatment effect assumption. In a more granular fashion, examining the $(j,j)$ entry of $\bm{V}_{c}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ and $\bm{V}(\bm{\xi}_{i\cdot})(a,a^{\prime})$ gives us
@@ 3.2.1 Without covariates
Let $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ denote the covariance matrix of $\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})=(\widehat{\tau}_{1,\mathrm{I}}(a,a^{\prime}),\ldots,\widehat{\tau}_{J,\mathrm{I}}(a,a^{\prime}))^{\top}$ and $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ denote the marginal variance of $\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})$, where $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ is the $(j,j)$ entry of $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$. Both $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ and $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ can be estimated by the CR variance estimator (Liang and Zeger, 1986), denoted by $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ and $\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$, respectively. For a WLS fit of (6), let $\bm{D}_{i,\mathrm{I}}$ and $\bm{\Pi}_{i,\mathrm{I}}$ respectively denote the design matrix and diagonal weight matrix for cluster $i$. We first obtain the following CR variance estimator for the vector of the regression coefficients,
$\bm{E}_{\mathrm{I}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{\top}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{I}}^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}\right)^{-1},$
@@
where $\widehat{\bm{e}}_{i,\mathrm{I}}$ is the $N_{i\cdot}$-dimensional vector of residuals from the WLS fit of (6), with $N_{i\cdot}=\sum_{j=1}^{J}N_{ij}$, $\bm{D}_{i,\mathrm{I}}$ is of dimensions $N_{i\cdot}\times J(J+1)$ with $\bm{D}_{i,\mathrm{I}}=(\bm{D}_{i1,\mathrm{I}}^{\top}\cdots\bm{D}_{iJ,\mathrm{I}}^{\top})^{\top}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}$ being $\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ijk}^{\top}$, and the $j$-th diagonal block of $\bm{\Pi}_{i,\mathrm{I}}$ is $\bm{\Pi}_{ij,\mathrm{I}}=\mathrm{diag}(\pi_{ij1},\ldots,\pi_{ijN_{ij}})$. Here, $\bm{G}_{i}(\mathcal{A})=(G_{i}(1),\ldots,G_{i}(J),G_{i}(\infty))^{\top}$, $\boldsymbol{\mathbbm{1}}_{ijk}=(\mathbbm{1}_{ijk,1}\ldots,\mathbbm{1}_{ijk,J})^{\top}$ with $\mathbbm{1}_{ijk,j^{\prime}}=1$ if $j=j^{\prime}$ and $\mathbbm{1}_{ijk,j^{\prime}}=0$ otherwise for $j^{\prime}=1,\ldots,J$, and $\otimes$’ denotes the Kronecker product such that $\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ijk}^{\top}=(G_{i}(1)\mathbbm{1}_{ijk}^{\top},\ldots,G_{i}(J)\mathbbm{1}_{ijk}^{\top},\allowbreak G_{i}(\infty)\mathbbm{1}_{ijk}^{\top})$. We can write $\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})=\bm{Q}(a,a^{\prime})\widehat{\bm{\beta}}_{\mathrm{I}}$, where $\widehat{\bm{\beta}}_{\mathrm{I}}=(\widehat{\bm{\beta}}_{\mathrm{I}}(1)^{\top},\ldots,\widehat{\bm{\beta}}_{\mathrm{I}}(J)^{\top},\widehat{\bm{\beta}}_{\mathrm{I}}(\infty)^{\top})^{\top}$, with $\widehat{\bm{\beta}}_{\mathrm{I}}(a)=(\widehat{\beta}_{1,\mathrm{I}}(a),\ldots,\widehat{\beta}_{J,\mathrm{I}}(a))^{\top}$ for $a\in\mathcal{A}$, and $\bm{Q}(a,a^{\prime})$ is a $J\times J(J+1)$-dimensional known coefficient matrix, e.g.,